# Graph Agent Network: Empowering Nodes with Inference Capabilities for Adversarial Resilience

### v2

@@ 1 Introduction
  • We experimentally show that primary functions of GAgN have been effectively executed, which collectively yielding state-of-the-art classification accuracy in perturbed datasets.
@@ 2 The Proposed Method
We start with action spaces and discuss the state’s structure and update strategies alongside other components to clarify the entire process. Figure 2 shows the workflow of GAgN.
@@ 2.6 Theoretical Analysis
This section provides the theoretical basis for designing inference function structures in agents, proving their computational universality for target tasks. We first establish the global equivalence between GAgN and GraphSAGE.
@@
Proof in Appendix A.1. We then propose the corollary to guide the design of the embedding function:
@@
Proof in Appendix A.2. For the structure of the degree inference function, we arrive at the following conclusions:
@@
Proof in Appendix A.3. For the neighboring confidence function’s structure, we present the following findings:
@@
$r_{p}$ RGCN SVD Pro Jaccard EGNN GAgN RGCN SVD Pro Jaccard EGNN GAgN
@@ 4 Conclusions
In this paper, we proposed GAgN, a novel and effectiveness approach for addressing the inherent vulnerabilities of GNNs to adversarial edge-perturbing attacks. By adopting a decentralized interaction mechanism, GAgN facilitates the filtration of adversarial edges and effectively thwarts global attack strategies. The theoretical sufficiency for GAgN further simplifies the model, while experimental results validate its effectiveness in resisting edge-perturbing attacks compared to existing defenses.
@@ A.3 Proof of Theorem 2
$\left(\mathbf{Z}\mathbf{Z}^{\top}+\alpha\Upsilon\right)^{-1}$ $=\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}-\alpha\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\Upsilon\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}$ (30)
@@
$\mathbf{Z}\mathbf{W}$ $=\mathbf{Z}\lim_{\alpha\to 0}\left(\left(\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O\left(\alpha\right)\right)\right)\mathbf{L}$ (32)
$=\lim_{\alpha\to 0}\left(\left(\mathbf{Z}\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\mathbf{L}+\mathbf{Z}O(\alpha)\mathbf{L}\right)\right)$ (33)

### v5

@@ 1 Introduction
  • We experimentally show that primary functions of GAgN have been effectively executed, which collectively yielding state-of-the-art accuracy in perturbed datasets.
@@ 2 The Proposed Method
We start with action spaces and discuss the state’s structure and update strategies alongside other components. Figure 2 shows the workflow of GAgN.
@@ 2.6 Theoretical Analysis
We first establish the global equivalence.
@@
We then propose the corollary to guide the design of the embedding function:
@@
For the structure of the degree inference function, we arrive at the following conclusions:
@@
For the neighboring confidence function’s structure, we present the following findings:
@@
Dataset $r_{p}$ RGCN SVD Pro Jaccard EGNN GAgN RGCN SVD Pro Jaccard EGNN GAgN
@@ 4 Conclusions
We proposed GAgN for addressing the inherent vulnerabilities of GNNs to edge-perturbing attacks. By adopting a decentralized interaction mechanism, GAgN facilitates the filtration of adversarial edges and thwarts global attacks. The theoretical sufficiency for GAgN further simplifies the model, while experimental results validate its effectiveness in resisting edge-perturbing attacks.
5 Acknowledgments
This work is supported in part by the National Key R&D Program of China (No. 2022YFB3102100, 2020YFB1805400), the National Science Foundation of China (No. U22B2027, 62032002, 62372313, 62172297, 62476107), the China Postdoctoral Science Foundation (No. 2024M752211), the Sichuan Science and Technology Program (No. 24QYCX0417), and the Youth Science Foundation of Sichuan (No. 25QNJJ4560).
@@ A.3 Proof of Theorem 2
$\left(\mathbf{Z}\mathbf{Z}^{\top}+\alpha\Upsilon\right)^{-1}$ $=\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}-\alpha\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\Upsilon\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}$
@@
$\mathbf{Z}\mathbf{W}$ $=\mathbf{Z}\lim_{\alpha\to 0}\left(\left(\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}+O\left(\alpha\right)\right)\right)\mathbf{L}$
$=\lim_{\alpha\to 0}\left(\left(\mathbf{Z}\mathbf{Z}^{\top}\left(\mathbf{Z}\mathbf{Z}^{\top}\right)^{-1}\mathbf{L}+\mathbf{Z}O(\alpha)\mathbf{L}\right)\right)$